b1az34:

Apple weights in an orchard are normally distributed. From a sample farmer, Fred determines the mean weight of a box of apples to be 270 oz. with a standard deviation of 10 oz. He wonders what percent of the apple boxes he has grouped for sale will have a weight of less than 255 oz.

b1az34:

1 attachment
b1az34:

A delivery driver has an average daily gasoline expense of $55.00. The standard deviation is $10.00. The owner takes a sample of 54 bills. What is the probability the mean of his sample will be between $45.00 and $65.00? Enter all answers to the nearest tenth. Calculate a z-score for $45.00. Give the probability for step 1. Calculate the z-score for $65.00. Give the probability for step 3. Add the probabilities from steps 1 & 3.

b1az34:

@addison123456 wrote:
i think it might be 2.07
for all

addison123456:

for the first question \

addison123456:

do you need help with the second one

b1az34:

yes

b1az34:

for both questions

addison123456:

k so When x =45, z 45-55 divided by 10 = -1 Step : z score = -1 Step 2: P(-1<Z<0) =0.3413 Step 3 : Z score for 65 = 65-55 divided by 10 = 1 Step 4: P(0<Z<1) = 0.3413 step 5: P(45<x<55) = 0.6826 answer is 0.6826 im pretty sure

addison123456:

thnx for da metal

b1az34:

@b1az34 wrote:
and this?

addison123456:

one second i didnt realize that was there lol

b1az34:

u still working it out?

addison123456:

ye

b1az34:

alr i'll wait

b1az34:

help

b1az34:

@addison123456 wrote:
ye
i heard you have a lil crush on me

TheyLuhhJae:

1 attachment
TheyLuhhJae:

1 attachment
TheyLuhhJae:

1 attachment
TheyLuhhJae:

1 attachment
TheyLuhhJae:

1 attachment
TheyLuhhJae:

1 attachment
TheyLuhhJae:

1 attachment
TheyLuhhJae:

1 attachment
TheyLuhhJae:

1 attachment
TheyLuhhJae:

1 attachment
TheyLuhhJae:

1 attachment
TheyLuhhJae:

1 attachment
TheyLuhhJae:

1 attachment
TheyLuhhJae:

1 attachment
TheyLuhhJae:

1 attachment
TheyLuhhJae:

1 attachment
TheyLuhhJae:

1 attachment
TheyLuhhJae:

1 attachment
TheyLuhhJae:

1 attachment
TheyLuhhJae:

1 attachment
TheyLuhhJae:

1 attachment
TheyLuhhJae:

1 attachment
TheyLuhhJae:

1 attachment
TheyLuhhJae:

1 attachment
TheyLuhhJae:

1 attachment